# Diff-Shadow: Global-guided Diffusion Model for Shadow Removal

### v1

@@ Abstract
We propose Diff-Shadow, a global-guided diffusion model for high-quality shadow removal. Previous transformer-based approaches can utilize global information to relate shadow and non-shadow regions but are limited in their synthesis ability and recover images with obvious boundaries. In contrast, diffusion-based methods can generate better content but ignore global information, resulting in inconsistent illumination. In this work, we combine the advantages of diffusion models and global guidance to realize shadow-free restoration. Specifically, we propose a parallel UNets architecture: 1) the local branch performs the patch-based noise estimation in the diffusion process, and 2) the global branch recovers the low-resolution shadow-free images. A Reweight Cross Attention (RCA) module is designed to integrate global contextural information of non-shadow regions into the local branch. We further design a Global-guided Sampling Strategy (GSS) that mitigates patch boundary issues and ensures consistent illumination across shaded and unshaded regions in the recovered image. Comprehensive experiments on three publicly standard datasets ISTD, ISTD+, and SRD have demonstrated the effectiveness of Diff-Shadow. Compared to state-of-the-art methods, our method achieves a significant improvement in terms of PSNR, increasing from 32.33dB to 33.69dB on the SRD dataset. Codes will be released.
1 Introduction
@@
Shadows are a phenomenon that exists when an optical image is captured with light blocked. The presence of shadows can complicate image processing as well as many subsequent vision tasks, e.g., object detection, tracking, and semantic segmentation [14]. The goal of shadow removal is to enhance the brightness of the image shadow regions and to achieve a consistent illumination distribution between shadow and non-shadow regions. The shadow removal task is essentially an image restoration task designed to utilize the information available in the shadow-affected image to recover the information lost due to light occlusion.
@@
Figure 1: (a) shows the result of ShadowFormer [14], which suffers from residual shadow artifacts due to the limited modeling ability, while the result of (b) ShadowDiffusion [15] shows obvious illumination inconsistency across the images because it cannot exploit the global information. (c) exhibits the proposed Diff-Shadow, which ultimately generates high-quality shadow removal results that maintain illumination consistency and are free from block boundary artifacts through the design of a parallel network structure and a novel global-guided sampling strategy.
@@
Shadow removal has received intensive and long-lasting studies. Traditional approaches use hand-crafted features and are based on physical lighting models [11, 47, 44]. Inspired by the successes of deep learning, many approaches that utilize large amounts of data [14, 2, 9, 49] have been introduced to improve shadow removal and reveal promising improvements. For example, Guo et al. proposed a transformer-based architecture, named ShadowFormer, to exploit non-shadow information to help shadow region restoration [14]. Their shadow-interaction module plays an important role in exploring global information. However, this kind of method often overlooks the modeling of the underlying distribution of natural images. Consequently, the generated results appear significant residual shadows, as shown in Fig. 1 (a).
@@
Recently, diffusion models have gained wide interest in the field of image restoration [33, 34, 19] due to its powerful modeling ability. Guo et al. proposed ShadowDiffusion to jointly pursue shadow-free images and refined shadow masks [15]. Although it covers the whole input image, the simple UNet denoising architecture is not enough for exploring global information, which results in inconsistent illumination of the outputs, as shown in Fig. 1 (b). Furthermore, generating arbitrary resolution images directly with ShadowDiffusion proved challenging as diffusion models are more suitable for processing smaller patches [6]. While Zheng et al. [6] and Ozan et al. [30] proposed several patch-based diffusion methods that demonstrated exceptional performance in image restoration tasks at arbitrary resolutions, they still faced limitations when handling shadow removal task. These methods lack global contextual information and ultimately lead to lighting inconsistencies or obvious boundaries in the restored images.
@@
Considering the powerful modeling capabilities of diffusion models and the necessity of global information for shadow removal tasks, we propose Diff-Shadow, a global-guided diffusion model for generating high-quality shadow-free results with no obvious boundaries and consistent illumination in arbitrary image resolution. Our Diff-Shadow adopts the parallel UNets network architecture that receives both patch and global inputs, where the local branch focuses on estimating patch noise, while the global branch is responsible for recovering low-resolution shadow-free images. Additionally, our Diff-Shadow exploits the global contextual information of the non-shadow region from the global branch through the Reweight Cross Attention (RCA) module. Subsequently, we present a Global-guided Sampling Strategy (GSS) within the denoising process. In this strategy, the fusion weights for patch noise are determined by considering both the brightness disparity between the recovered patch and the global image, as well as the extent of shadow regions within the patch. This approach addresses potential merging artifacts, preserves excellent illumination consistency, and ultimately enhances the overall quality of the results, as shown in Fig. 1 (c). Moreover, benefiting from the patch-based methods, the proposed Diff-Shadow can be expanded to handle images with arbitrary resolution.
@@
  • We propose Diff-Shadow, a global-guided Diffusion Model pipeline, which incorporates a parallel UNets architecture, concentrating on both the patch and global information for the effective shadow removal.
  • We design a Reweight Cross Attention (RCA) module to exploit the correlation between the local branch and the non-shadow regions of the global image.
@@
  • We propose a Global-guided Sampling Strategy (GSS) to further improve the quality of shadow removal during the diffusion denoising process.
@@
  • Extensive experimental results on public ISTD [41], ISTD+ [22], and SRD [31] datasets demonstrate that the proposed Diff-Shadow outperforms the state-of-the-art shadow removal methods.
2 Related Works
Shadow Removal.
@@ 2 Related Works
Recent learning-based methods for shadow removal tasks have achieved superior performance by leveraging large-scale high-quality training data to learn high-level image texture and semantic information [2, 9, 14, 15]. Unlike traditional physical-based methods that rely on prior assumptions like illumination [1, 47], gradient [7, 8, 27, 25, 11], and region [16, 12, 13], learning-based approaches tend to learn contextual features for shadow removal. For instance, Qu et al. proposed a multi-context embedding network to integrate multi-level information [31], while Wang et al. explored the relationship between shadow detection and removal, and proposed a stacked conditional generative adversarial network (ST-CGAN) [41]. Other works like Hu et al. [18] and Zhang et al. [48] focused on applying spatial context attention features and exploring residual-illumination relationships, respectively. These learning-based methods have produced impressive results, but there remain opportunities to further capture the correspondence between shadow and non-shadow areas.
Diffusion Methods.
Recent advances in generative diffusion models [17] , [28] and noise-conditional score networks  [36][37] have shown powerful image synthesis capabilities [5]. By gradually corrupting images into noise and reversing the process, high-fidelity and diverse results can be sampled. Therefore, many studies have been inspired by the powerful generative capability of diffusion models and applied them to image restoration tasks, such as super-resolution [34, 10, 29, 24, 35, 32], deblurring [43, 3], denoising [45, 38], and inpainting [33, 46, 26]. These tasks have achieved promising results, with some even reaching the state-of-the-art level. For general image restoration, Özdenizci et al. proposed a patch-based denoising diffusion approach to handle arbitrary image sizes [30]. By operating on image patches, their method restored local details while maintaining global coherence via overlapping patches. As for the shadow removal task, ShadowDiffusion integrates image and degradation priors into a unified diffusion framework [15], which progressively refines the output image and the shadow mask for accurate shadow-free image generation. We design a diffusion model with parallel UNets architecture to perform the local patch noise estimation and the global image reconstruction.
@@
Figure 2: An overview of the proposed Diff-Shadow. The local branch of Parallel UNets performs patch diffusion noise estimation using patches of the intermediate variable $\mathbf{x}_{t}$, the shadow image $\tilde{\mathbf{x}}$, and the shadow mask $\mathbf{x}_{m}$. The latter two images are also down-sampled as inputs of the global UNet, which constructs the low-resolution shadow-free image on the one hand and provides the global contextual information of non-shaded regions $\mathbf{x}_{G}$ into the local branch ${\mathbf{x}_{L}}^{(i)}$ using the proposed Reweight Cross Attention (RCA) module on the other hand. After estimating the noises of local patches, the Global-guided Sampling Strategy (GSS) works to construct the noise distribution for the whole image. $t_{1}$ and $t_{2}$ represent functions corresponding to the step $t$.
Shadow and Non-Shadow Consistency.
@@
Shadow removal presents a unique challenge in that shadows corrupt only parts of an image, while abundant uncorrupted information remains in non-shadowed regions. However, previous methods [22, 23] focused only on local shadow or non-shadow regions, leading to severe artifacts around shadow boundaries and inconsistent brightness between shadow and non-shadow areas. More recent approaches leverage contextual information from both shadow and non-shadow areas of the image to inform the shadow removal process. For instance, DeshadowNet [31] incorporates multi-scale information to expand the receptive field, leveraging texture and appearance information across the image. Chen et al. proposed CANet [2], using an external patch-matching module to discover texture relationships. Guo et al. introduced ShadowFormer [14], incorporating a Shadow-Interaction Module in the transformer to model the global texture. Related but divergent, we propose the Reweight Cross Attention module injected into the diffusion model, empowering it with global textural relationship modeling.
3 Method
@@
In this section, we describe the proposed global-guided denoising diffusion model for shadow removal (Diff-Shadow). Given the intermediate variable $\mathbf{x}_{\text{t}}\in\mathbb{R}^{3\times H\times W}$, the shadow image $\tilde{\mathbf{x}}\in\mathbb{R}^{3\times H\times W}$ with corresponding shadow mask $\mathbf{x}_{m}\in\mathbb{R}^{3\times H\times W}$ in Fig. 2 (a), the proposed architecture is trained and inference using both high-resolution image patches as well as low-resolution down-sampled full images, which are optimized by the parallel UNets (Fig. 2 (d)) whose local branch performs the diffusion noise estimation of local patches and the global UNet recovers the down-sampled full shadow-free image. The global contextual information of non-shadow regions is integrated into the local branch through the RCA module (Fig. 2 (f)) and the Convolutional Aggregation (CA) during training. The Global-guided Sampling Strategy (GSS) (Fig. 2 (e)) is utilized to systematically combine the features of neighboring patches during inference. Thanks to the excellent modeling ability of the diffusion model with small-size images, the proposed Diff-Shadow achieves satisfactory performance for shadow removal tasks.
3.1 Overall Architecture
@@
Condition diffusion model [34] have achieved impressive performance for image-conditional data synthesis and editing tasks, which learns a conditional reverse process $p_{\theta}\left(\mathbf{x}_{0:T}\mid\tilde{\mathbf{x}}\right)$ without modifying the diffusion process $q\left(\mathbf{x}_{1:T}\mid\mathbf{x}_{0}\right)$ for ${\mathbf{x}}$, such that the sampled image has high fidelity to the data distribution conditioned on $\tilde{\mathbf{x}}$. Common diffusive models train samples $\left(\mathbf{x}_{0},\tilde{\mathbf{x}}\right)\sim q\left(\mathbf{x}_{0},\tilde{\mathbf{x}}\right)$ from a paired data distributions (e.g., a clean image $\mathbf{x}_{0}$ and a degraded image $\tilde{\mathbf{x}}$), and learn a conditional diffusion model where it provides $\tilde{\mathbf{x}}$ to the reverse process:
$\small p_{\theta}\left(\mathbf{x}_{0:T}\mid\tilde{\mathbf{x}}\right)=p\left(\mathbf{x}_{T}\right)\prod_{t=1}^{T}p_{\theta}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},\tilde{\mathbf{x}}\right).$ (1)
@@
As for shadow removal tasks with shadow masks $\mathbf{x}_{m}$, we design the parallel UNets architecture. The local branch performs the diffusion noise estimation and is conditioned by image patches: ${\mathbf{x}_{L}}^{(i)}=\left\{\tilde{{\mathbf{x}}}^{(i)},{\mathbf{x}_{m}}^{(i)}\right\}$, where $\tilde{\mathbf{x}}^{(i)}=\operatorname{Crop}\left(\boldsymbol{L}_{i}\circ\tilde{\mathbf{x}}\right)$ and ${\mathbf{x}_{m}}^{(i)}=\text{Crop}\left(\boldsymbol{L}_{i}\circ\mathbf{x}_{m}\right)$ represent $R\times R$ patches, and $\text{Crop}(\cdot)$ extracts the patches from the location $\boldsymbol{L}_{i}$. The estimated noise of local patches will be combined during the inference through the proposed GSS modules (Fig. 2 (e)), which is described in Sec. 3.3. The global branch is optimized by the low-resolution down-sampled full images: $\mathbf{x}_{G}=\left\{\tilde{{\mathbf{x}}}^{\prime},{\mathbf{x}_{m}}^{\prime}\right\}$. The contextual information of non-shadow regions plays an auxiliary role and is integrated into the local branch through the RCA module (Fig. 2 (f)), which is described in Sec. 3.2. Therefore, the reverse process in Eq. 1 can be expanded as:
$\footnotesize p_{\theta}\!\left(\mathbf{x}_{0:T}^{(i)}\!\mid\!{\mathbf{x}_{L}}^{(i)},\mathbf{x}_{G}\right)\!=\!p\left(\mathbf{x}_{T}^{(i)}\right)\prod_{t=1}^{T}p_{\theta}\!\left(\mathbf{x}_{t-1}^{(i)}\!\mid\!\mathbf{x}_{t}^{(i)},{\mathbf{x}_{L}}^{(i)},\mathbf{x}_{G}\right),$ (2)
@@
$\small\boldsymbol{{\epsilon}_{t}}=\boldsymbol{\epsilon}_{\theta}\left(\sqrt{\bar{\alpha}_{t}}{\mathbf{x}_{0}}^{(i)}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon},{\mathbf{x}_{L}}^{(i)},\mathbf{x}_{G},t\right).$ (3)
@@
3.2 Reweight Cross Attention
Existing patch-based approaches lack generality since the regions inside a patch may all be corrupted for shadow removal tasks. Apart from the local patch-based diffusion branch, we introduce a parallel global UNet to recover low-resolution shadow-free images. The Reweight Cross Attention (RCA) module is introduced to integrate the global information into the local patch by combining the low-level features. As shown in Fig. 2 (f), the output of RCA can be summarized as:
$Z=\operatorname{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{d}\right)\mathbf{M}\mathbf{V},$ (4)
where $\mathbf{Q}$ represents the projected queries of patch features, $\mathbf{K}$ and $\mathbf{V}$ are the projected keys and values of global image features, $d$ is the scaling parameter, and $\mathbf{M}$ is the binary mask to multiply with the features. Attention maps between $\mathbf{Q}$ and $\mathbf{K}$ are first calculated to capture mutual information between pixels in the local branch and regions in the global branch. Subsequently, the mask is introduced to assist in suppressing mutual information values in the global shadow regions. The adjusted attention map is then multiplied by the global $\mathbf{V}$ vector to obtain the expected global information for the local branch.
We interact the low-level features of the first convolutional block in the global UNet with the corresponding layer in the local UNet through the RCA module because low-level features concentrate more on image edges, colors, and corners, which are important for shadow removal tasks. As for high-level abstract features, we apply a simple concatenation to achieve interaction. In contrast to recent cross-attention modules in Shadowformer [14] whose $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$ originate from the same global feature, the RCA integrates both local and global information. Furthermore, other attention-based methods boost the mutual information between shadow and non-shadow pairs through masks. However, the shadow information is dispensable for recovering non-shadow regions. Contrarily, our attention mechanism utilizes masks to suppress the global shadow values through the attention map, integrating the more important non-shadow information from the global branch into the local branch.
@@ 3.2 Reweight Cross Attention
3.3 Global-guided Samping Strategy
To overcome the shortcomings of merging artifacts from independently intermediate patches, Ozdenizci et al. proposed to perform reverse sampling based on the mean estimated noise for each pixel in overlapping patch regions, which works better for weather image with relatively unified degradation [30]. However, as for shadow removal tasks with extreme inconsistencies between shadow and non-shadow regions, predicting mean estimated noise lacks ability to achieve a consistent illumination distribution. Therefore, we propose the Global-guided Sampling Strategy (GSS) to preserve excellent illumination consistency.
The general idea of patch-based restoration is to operate locally on patches extracted from the image and optimally merge the results.
We consider a grid-like arranged parsing scheme over the intermediate variable $\mathbf{x}_{t}$, the shadow image $\tilde{\mathbf{x}}$ and the shadow mask $\mathbf{x}_{m}$, and extract all patches by moving over the $R\times R$ grid with a step size of $r$ in both horizontal and vertical dimensions, as shown in Fig. 3. During patch-based diffusive image restoration, we perform noise estimation based on adaptive estimated noise for each pixel in overlapping patch regions, at any given denoising time step $t$. As shown in Fig. 2 (e), the weight of each patch is determined by two parts: the brightness-based score $S_{\text{bri}}$ and the mask-based score $S_{\text{m}}$. The former one considers the brightness consistency of the local patches and the global images, while the latter one corresponds to the proportion of shadow and non-shadow regions.
@@ 3.3 Global-guided Samping Strategy
$\small S_{\text{bri}}=1-\frac{\mathrm{avg}{(|L_{\text{gray}}-G_{\text{gray}}|)}}{\mathrm{max}({(|L_{\text{gray}}-G_{\text{gray}}|)})+\mathrm{eps}},$ (5)
@@
$\small S_{\text{m}}=\left\{\begin{array}[]{cc}1&R_{\text{sf}}\geq T_{\text{sf}}\\ R_{\text{sf}}/{T_{\text{sf}}}&R_{\text{sf}}<T_{\text{sf}}\end{array},\right.$ (6)
@@
$\small W^{d}=\frac{t\cdot S_{\text{bri}}}{T}+\frac{(T-t)\cdot S_{\text{m}}}{T},$ (7)

### v2

@@ Abstract
We propose Diff-Shadow, a global-guided diffusion model for shadow removal. Previous transformer-based approaches can utilize global information to relate shadow and non-shadow regions but are limited in their synthesis ability and recover images with obvious boundaries. In contrast, diffusion-based methods can generate better content but they are not exempt from issues related to inconsistent illumination. In this work, we combine the advantages of diffusion models and global guidance to achieve shadow-free restoration. Specifically, we propose a parallel UNets architecture: 1) the local branch performs the patch-based noise estimation in the diffusion process, and 2) the global branch recovers the low-resolution shadow-free images. A Reweight Cross Attention (RCA) module is designed to integrate global contextual information of non-shadow regions into the local branch. We further design a Global-guided Sampling Strategy (GSS) that mitigates patch boundary issues and ensures consistent illumination across shaded and unshaded regions in the recovered image. Comprehensive experiments on datasets ISTD, ISTD+, and SRD have demonstrated the effectiveness of Diff-Shadow. Compared to state-of-the-art methods, our method achieves a significant improvement in terms of PSNR, increasing from 32.33dB to 33.69dB on the ISTD dataset.
Codehttps://github.com/MonteCarluo/Diff-Shadow
Introduction
@@
Shadows are a phenomenon that exists when an optical image is captured with light blocked. The presence of shadow can complicate image processing as well as many subsequent vision tasks, e.g., object detection, tracking, and semantic segmentation (Guo et al. 2023a). The goal of shadow removal is to enhance the brightness of the image shadow regions and to achieve a consistent illumination distribution between shadow and non-shadow regions. The shadow removal task is essentially an image restoration task designed to utilize the information available in the shadow-affected image to recover the information lost due to light occlusion.
@@
Figure 1: (a) shows the result of ShadowFormer (Guo et al. 2023a), which suffers from residual shadow artifacts due to the limited modeling ability, while the result of (b) ShadowDiffusion (Guo et al. 2023b) shows obvious illumination inconsistency across the images because it cannot exploit the global information. (c) exhibits the proposed Diff-Shadow, which generates high-quality shadow removal results that maintain illumination consistency and are free from block boundary artifacts through the design of a parallel network structure and a novel global-guided sampling strategy.
@@
Shadow removal has received intensive and long-lasting attention. Traditional approaches use hand-crafted features and are based on physical lighting models (Gryka, Terry, and Brostow 2015; Zhang, Zhang, and Xiao 2015; Xiao et al. 2013). Inspired by the successes of deep learning, many approaches that utilize large amounts of data (Guo et al. 2023a; Chen et al. 2021; Fu et al. 2021; Zhu et al. 2022a) have been introduced to improve shadow removal and reveal promising improvements,  e.g., Guo et al. proposed a transformer-based architecture, named ShadowFormer, to exploit non-shadow information to help shadow region restoration. Their shadow-interaction module plays an important role in exploring global information. However, this method often overlooks the modeling of the underlying distribution of natural images. Consequently, the generated results appear to have significant residual shadows, as shown in Fig. 1 (a).
Recently, diffusion models have gained wide interest in the field of image restoration (Saharia et al. 2022a, b; Jiang et al. 2023) due to its powerful modeling ability. Guo et al. proposed ShadowDiffusion to jointly pursue shadow-free images and refined shadow masks (Guo et al. 2023b).
@@
Although it covers the whole input image, the kind of conventional diffusion-based methods use a simple UNet and are challenging for shadow regions to capture effective information from matched non-shadow regions, resulting in illumination inconsistency in Fig. 1 (b). Furthermore, generating arbitrary resolution images directly with ShadowDiffusion proved challenging as diffusion models are more suitable for processing smaller patches (Ding et al. 2023). While Zheng et al. (Ding et al. 2023) and Ozan et al. (Özdenizci and Legenstein 2023) proposed several patch-based diffusion methods that demonstrated exceptional performance in image restoration tasks at arbitrary resolutions, they still faced limitations when handling shadow removal task. These methods lack global contextual information and ultimately lead to lighting inconsistencies or obvious boundaries in the restored images.
@@
Considering the powerful modeling capabilities of diffusion models and the necessity of global information for shadow removal tasks, we propose Diff-Shadow, a global-guided diffusion model for generating high-quality shadow-free results with no obvious boundaries and consistent illumination in arbitrary image resolution. Our Diff-Shadow adopts the parallel UNets network architecture that receives both patch and global inputs, where the local branch focuses on estimating patch noise, while the global branch is responsible for recovering low-resolution shadow-free images. Additionally, our Diff-Shadow exploits the global contextual information of the non-shadow region from the global branch through the Reweight Cross Attention (RCA) module. Subsequently, we present a Global-guided Sampling Strategy (GSS) within the denoising process. In this strategy, the fusion weights for patch noise are determined by considering both the brightness disparity between the recovered patch and the global image, as well as the extent of shadow regions within the patch. This approach addresses potential merging artifacts, preserves superb illumination coherence across shaded and non-shadow regions, and ultimately enhances the overall quality of the results, as shown in Fig. 1 (c). Moreover, benefiting from the patch-based methods, the Diff-Shadow can be expanded to handle images with arbitrary resolution.
@@
(1) We propose Diff-Shadow, a global-guided diffusion architecture that incorporates parallel UNets and exploits the correlation between the local branch and the non-shadow regions of the global through a reweighted cross-attention (RCA) module, which removes shadows efficiently.
@@
(2) We propose a Global-guided Sampling Strategy (GSS) to further improve the quality of shadow removal during the diffusion denoising process.
@@
(3) Extensive experimental results on public ISTD (Wang, Li, and Yang 2018), ISTD+ (Le and Samaras 2019), and SRD (Qu et al. 2017) datasets demonstrate that the proposed Diff-Shadow outperforms the state of the art.
Related Works
@@
Shadow Removal. Recent learning-based methods for shadow removal tasks have achieved superior performance by leveraging large-scale high-quality training data to learn high-level image texture and semantic information (Chen et al. 2021; Fu et al. 2021; Guo et al. 2023a, b). Unlike traditional physical-based methods that rely on prior assumptions like illumination (Arbel and Hel-Or 2010; Zhang, Zhang, and Xiao 2015), gradient (Finlayson, Hordley, and Drew 2002; Finlayson et al. 2005; Mohan, Tumblin, and Choudhury 2007; Liu and Gleicher 2008; Gryka, Terry, and Brostow 2015), and region (Guo, Dai, and Hoiem 2012; Guo et al. 2021, 2022), the learning-based approaches tend to learn contextual features for shadow removal. For instance, Qu et al. proposed a multi-context embedding network to integrate multi-level information (Qu et al. 2017), while Wang et al. explored the relationship between shadow detection and removal, and proposed a stacked conditional generative adversarial network (ST-CGAN) (Wang, Li, and Yang 2018). Zhu et al. proposed a model-driven network (Zhu et al. 2022b) combining efficiency and interpretability by formulating a new shadow illumination model and unfolding it into a deep network. Other works like Hu et al. (Hu et al. 2019) and Zhang et al. (Zhang et al. 2020) focused on applying spatial context attention features and exploring residual-illumination relationships, respectively. These learning-based methods have produced impressive results, but there remain opportunities to further capture the correspondence between shadow and non-shadow areas.
@@
Figure 2: An overview of the Diff-Shadow. The local branch of Parallel UNets performs patch diffusion noise estimation using patches of the intermediate variable $\mathbf{x}_{t}$, the shadow image $\tilde{\mathbf{x}}$, and the shadow mask $\mathbf{x}_{m}$. The latter two images are also down-sampled as inputs of the global UNet, which constructs the low-resolution shadow-free image on the one hand and provides the global contextual information of non-shaded regions $\mathbf{x}_{G}$ into the local branch ${\mathbf{x}_{L}}^{(i)}$ using the proposed Reweight Cross Attention (RCA) module on the other hand. After estimating the noises of local patches, the Global-guided Sampling Strategy (GSS) works to construct the noise distribution for the whole image. $t_{1}$ and $t_{2}$ represent functions corresponding to the step $t$.
@@
Shadow and Non-Shadow Consistency. Shadow removal presents a unique challenge in that shadows corrupt only parts of an image, while abundant uncorrupted information remains in non-shadowed regions. However, previous methods (Le and Samaras 2019, 2020) focused only on local shadow or non-shadow regions, leading to severe artifacts around shadow boundaries and inconsistent brightness between shadow and non-shadow areas. More recent approaches leverage contextual information from both shadow and non-shadow areas of the image to inform the shadow removal process. For instance, DeshadowNet (Qu et al. 2017) incorporates multi-scale information to expand the receptive field, leveraging texture and appearance information across the image. Chen et al. proposed CANet (Chen et al. 2021), using an external patch-matching module to discover texture relationships. Jin et al. proposed SG-ShadowNet (Wan et al. 2022a), reformulating shadow removal as an intra-image style transfer problem to harmonize de-shadowed regions with non-shadow areas. Guo et al. introduced ShadowFormer (Guo et al. 2023a), incorporating a Shadow-Interaction Module in the transformer to model the global texture. Related but divergent, we propose the Reweight Cross Attention module injected into the diffusion model, empowering it with global textural relationship modeling.
Method
@@
Given the intermediate variable $\mathbf{x}_{\text{t}}\in\mathbb{R}^{3\times H\times W}$, the shadow image $\tilde{\mathbf{x}}\in\mathbb{R}^{3\times H\times W}$ with corresponding shadow mask $\mathbf{x}_{m}\in\mathbb{R}^{1\times H\times W}$ in Fig. 2 (a), the proposed architecture is trained and inference using both high-resolution image patches as well as low-resolution down-sampled full images, which are optimized by the parallel UNets (Fig. 2 (d)) whose local branch performs the diffusion noise estimation of local patches and the global UNet recovers the down-sampled full shadow-free image. The global information of non-shadow regions is integrated into the local branch through the RCA module (Fig. 2 (f)) and the Convolutional Aggregation (CA) during training. The Global-guided Sampling Strategy (GSS) (Fig. 2 (e)) is utilized to combine the features of neighboring patches during inference.
Overall Architecture
@@
Condition diffusion model (Saharia et al. 2022b) have achieved impressive performance for image-conditional data synthesis and editing tasks, which learns a conditional reverse process $p_{\theta}\left(\mathbf{x}_{0:T}\mid\tilde{\mathbf{x}}\right)$ without modifying the diffusion process $q\left(\mathbf{x}_{1:T}\mid\mathbf{x}_{0}\right)$ for ${\mathbf{x}}$, such that the sampled image has high fidelity to the data distribution conditioned on $\tilde{\mathbf{x}}$. Common diffusive models train samples $\left(\mathbf{x}_{0},\tilde{\mathbf{x}}\right)\sim q\left(\mathbf{x}_{0},\tilde{\mathbf{x}}\right)$ from a paired data distributions ( e.g., a clean image $\mathbf{x}_{0}$ and a degraded image $\tilde{\mathbf{x}}$), and learn a conditional diffusion model that provides $\tilde{\mathbf{x}}$ to the reverse process:
$p_{\theta}\left(\mathbf{x}_{0:T}\mid\tilde{\mathbf{x}}\right)=p\left(\mathbf{x}_{T}\right)\prod_{t=1}^{T}p_{\theta}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},\tilde{\mathbf{x}}\right).$ (1)
@@
As for shadow removal tasks with shadow masks $\mathbf{x}_{m}$, we design the parallel UNets architecture. The local branch performs the diffusion noise estimation and is conditioned by image patches: ${\mathbf{x}_{L}}^{(i)}=\left\{\tilde{{\mathbf{x}}}^{(i)},{\mathbf{x}_{m}}^{(i)}\right\}$, where $\tilde{\mathbf{x}}^{(i)}=\operatorname{Crop}\left(\boldsymbol{L}_{i}\circ\tilde{\mathbf{x}}\right)$ and ${\mathbf{x}_{m}}^{(i)}=\text{Crop}\left(\boldsymbol{L}_{i}\circ\mathbf{x}_{m}\right)$ represent $R\times R$ patches, and $\text{Crop}(\cdot)$ extracts the patches from the location $\boldsymbol{L}_{i}$. The estimated noise of local patches will be combined during the inference through the proposed GSS module (Fig. 2 (e)). The global branch is optimized by the low-resolution down-sampled full images: $\mathbf{x}_{G}=\left\{\tilde{{\mathbf{x}}}^{\prime},{\mathbf{x}_{m}}^{\prime}\right\}$. The contextual information of non-shadow regions plays an auxiliary role and is integrated into the local branch through the RCA module (Fig. 2 (f)). Therefore, the reverse process in Eq. 1 can be expanded as:
$p_{\theta}\!\left(\mathbf{x}_{0:T}^{(i)}\!\mid\!{\mathbf{x}_{L}}^{(i)},\mathbf{x}_{G}\right)\!=\!p\left(\mathbf{x}_{T}^{(i)}\right)\prod_{t=1}^{T}p_{\theta}\!\left(\mathbf{x}_{t-1}^{(i)}\!\mid\!\mathbf{x}_{t}^{(i)},{\mathbf{x}_{L}}^{(i)},\mathbf{x}_{G}\right),$ (2)
@@
$\boldsymbol{{\epsilon}_{t}}=\boldsymbol{\epsilon}_{\theta}\left(\sqrt{\bar{\alpha}_{t}}{\mathbf{x}_{0}}^{(i)}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon},{\mathbf{x}_{L}}^{(i)},\mathbf{x}_{G},t\right).$ (3)
@@
Reweight Cross Attention
Existing patch-based methods lack generality since the regions inside a patch may all be corrupted. Apart from the local patch-based diffusion branch, we introduce a parallel global UNet to recover low-resolution shadow-free images. The Reweight Cross Attention (RCA) module is introduced to integrate the global information into the local patch by combining the low-level features. As shown in Fig. 2 (f), the output of RCA is:
$\mathbf{Z}=\operatorname{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{d}\right)\mathbf{M}\mathbf{V},$ (4)
$\mathbf{Q}$ represents the projected queries of patch features, $\mathbf{K}$ and $\mathbf{V}$ are the projected keys and values of global image features, $d$ is the scaling parameter, $\mathbf{M}$ is the binary mask to multiply with the features, and $\mathbf{Z}$ represents the output interacted features. Attention maps between $\mathbf{Q}$ and $\mathbf{K}$ are first calculated to capture the mutual information between pixels in the local branch and regions in the global branch. Subsequently, the mask is introduced to assist in suppressing mutual information values in the global shadow regions. The adjusted attention map is then multiplied by the global $\mathbf{V}$ vector to obtain the expected global information for the local branch.
We interact the low-level features of the first convolutional block in the global UNet with the corresponding layer in the local UNet through the RCA because low-level features concentrate more on image edges, colors, and corners, which are important for shadow removal tasks. As for high-level abstract features, we apply a simple concatenation to achieve interaction. In contrast to recent cross-attention modules in ShadowFormer (Guo et al. 2023a) whose $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$ originate from the same global feature, the RCA integrates both local and global information. Furthermore, other methods boost the mutual information between shadow and non-shadow pairs through masks. However, the shadow information is dispensable for recovering non-shadow regions. Contrarily, our attention mechanism utilizes masks to suppress the global shadow values through the attention map, integrating the more important non-shadow information from the global branch into the local branch.
@@
Global-guided Samping Strategy
To overcome the shortcomings of merging artifacts from independently intermediate patches, Ozdenizci et al. proposed to perform reverse sampling based on the mean estimated noise for each pixel in overlapping patch regions, which works better for weather image with relatively unified degradation (Özdenizci and Legenstein 2023). However, as for shadow removal tasks with extreme inconsistencies between shadow and non-shadow regions, predicting mean estimated noise lacks ability to achieve a consistent illumination distribution. Therefore, we propose the Global-guided Sampling Strategy (GSS) to maintain illumination consistency.
The general idea of patch-based restoration is to operate locally on patches extracted from the image and optimally merge the results. We consider a grid-like arranged parsing scheme over the intermediate variable $\mathbf{x}_{t}$, the shadow image $\tilde{\mathbf{x}}$ and the shadow mask $\mathbf{x}_{m}$, and extract all patches by moving over the $R\times R$ grid with a step size of $r$ in both horizontal and vertical dimensions, as shown in Fig. 3. During patch-based diffusive image restoration, we perform noise estimation based on adaptive estimated noise for each pixel in overlapping patch regions, at any given denoising time step $t$. As shown in Fig. 2 (e), the weight of each patch is determined by two parts: the brightness-based score $S_{\text{bri}}$ and the mask-based score $S_{\text{m}}$. The former one considers the brightness consistency of the local patches and the global images, while the latter one corresponds to the proportion of shadow and non-shadow regions.
@@
$S_{\text{bri}}=1-\frac{\mathrm{avg}{(|L_{\text{gray}}-G_{\text{gray}}|)}}{\mathrm{max}({(|L_{\text{gray}}-G_{\text{gray}}|)})},$ (5)
@@
$S_{\text{m}}=\left\{\begin{array}[]{cc}1&R_{\text{sf}}\geq T_{\text{sf}}\\ R_{\text{sf}}/{T_{\text{sf}}}&R_{\text{sf}}<T_{\text{sf}}\end{array},\right.$ (6)
@@
$W^{d}=\frac{t\cdot S_{\text{bri}}}{T}+\frac{(T-t)\cdot S_{\text{m}}}{T},$ (7)